# Transfer Matrix Approach for Topological Edge States

### v1

@@ Abstract
We suggest and develop a novel approach for describing topological properties of a periodic system purely from the transfer matrix associated to a unit cell. This approach is employed to obtain a simple pictorial proof of the Zak phase bulk-edge correspondence for any one dimensional system. We apply the transfer matrix method to study some generalisations of the Su-Schrieffer-Heeger (SSH) model, such as the tetramer SSH4 model and a dimerised one-dimensional photonic crystal.
@@ I Introduction
Of these topological insulators, one of the simplest toy models is the Su-Schrieffer-Heeger (SSH) model, which was first introduced to model electron hopping in long-chain polyenes [6]. Despite its simplicity, the SSH model and its generalisations are often used in experiment to demonstrate techniques for detecting topological phase transitions [7], as well as to demonstrate practical applications of topological edge states [8]. Theoretically, the SSH model and its generalisations serve as a starting point to test new methods for analysing topological insulators [9, 10]. Generalisations of the SSH model often also admit exciting nontrivial behaviour, which may shed light on the nature of topological phases [11], or present opportunities for technological applications [12]. On the other hand, due to its simplicity, many models in different fields reduce to the SSH model. For example, the SSH model has recently been realised in acoustic [13] and magnonic [14] systems.
The paper is organized as follows. In Sec. II, we begin by reviewing the transfer matrix method and apply it to solve the semi-infinite SSH model. In Sec. III, we demonstrate how this the transfer matrix approach can be employed to obtain topological invariants for the Hamiltonian. This method involves assigning to each Hamiltonian a path on a manifold with the endpoints fixed. We then show how our method connects to bulk-band topological invariants such as the Zak phase. In doing so, we obtain a short, but rigorous proof of the bulk-edge correspondence in general one dimensional systems possessing inversion symmetry. Finally, in Sec. IV we show how our method can be applied to more complicated models such as the generalized SSH model and a dimerised photonic crystal.
@@ II Transfer Matrix of the SSH model
The transfer ($T$) matrix is a powerful method of solving linear one-dimensional systems. It has previously also been considered in the study of topological insulators [15, 16, 17, 18], though our approach is considerably different. To derive the transfer matrix, it is convenient to rewrite the time-independent Schrödinger equation $H\psi=E\psi$ in matrix form:
@@
where $\bm{a}_{i}=(a_{2i},a_{2i+1})^{T}$ is the amplitudes in each unit cell. This process also shows that in general, any one-dimensional periodic dynamical system admits a unique set of transfer matrices.
The main advantage of the $T$-matrix formalism appears when we consider semi-infinite and finite chains. For the SSH model, the Dirichlet boundary conditions as applied in Eq. (2) is equivalent to setting the amplitude to the left of the first amplitude to be zero. As a result, the amplitudes over the entire chain are determined by the amplitude at the first site by the $T$-matrix:
@@
The only cases where Bloch’s theorem still applies are the cases when both eigenvalues of $T$ are equal, and edge modes, where the left and right edge modes do not couple. So, the left edge mode is obtained by the choice of $E$ which makes $(0,1)^{T}$ an eigenvector of the $T$ matrix. This is easily read off from (5) to be $\mu_{1}=v_{1}-E=0$. Finally, substituting this into our $T$-matrix, we recover the well-known condition that the left edge mode exists if and only if $|\gamma_{1}/\gamma_{2}|<1$. The main advantage of this technique is that we are able to obtain the condition for an edge mode in the SSH model without any symmetry or hermiticity requirement on the Hamiltonian. As such, the condition $|\gamma_{1}/\gamma_{2}|<1$ works just as well for $\gamma_{1},\gamma_{2}\in\mathbb{C}$ (non-hermitian models, e.g. systems with gain, nonreciprocal hopping, etc), as well as for symmetry-broken SSH models with jagged on-site potentials.
@@ III Topological Protection in the Transfer Matrix Formalism
Figure 1: (a) Illustration of the SSH model. The open circles denote sites at which the electron can sit, and the lines denote the coupling strength between each site. In the SSH model, the couplings alternate between two values $\gamma_{1}$ and $\gamma_{2}$. (b) The band structure of the SSH model $(\mu_{1}=\mu_{2})$ embedded in $\mathrm{SL}(2,\mathbb{R})$ (3D Iwasawa parameters space). Each curve represents a Hamiltonian, or equivalently, a family of $T$-matrices $T(E)$ for a different value of $\gamma_{1}/\gamma_{2}$. When the curve touches one of the orange surfaces, an edge state is supported. The decay length of this edge state is given by $|\ln(a)|^{-1}$. The green regions are the allowed bands: Each point on the curve inside the green region corresponds to a bulk propagating mode.
@@ III.1 The $\mathbb{R}$) Picture for Topological Phases
$T=\begin{pmatrix}\cos(k)&-\sin(k)\\ \sin(k)&\cos(k)\end{pmatrix}\begin{pmatrix}a&\\ &\frac{1}{a}\end{pmatrix}\begin{pmatrix}1&n\\ &1\end{pmatrix}$ (6)
where $k,n\in\mathbb{R},a>0$ are the Iwasawa parameters, and they are smooth functions of the entries $T_{ij}$. By doing so, we can parameterise $T(E)$ as a curve in a three-dimensional space, where the coordinates are taken to be the Iwasawa parameters: $k(E)$, $a(E)$, $n(E)$. We plot a family of curves with various choices of $\gamma$ in Fig. 1 for the SSH model.
@@ III.2 Proof of the Zak Phase Bulk-Edge Correspondence
Here, we show how our formalism so far is connected to Bloch band topological invariants. In doing so, we will obtain a rigorous but pictorially simple proof of the bulk-edge correspondence. We stress that our results hold for any number of sites per unit cell, as well as for continuous quantum mechanical or photonic systems.
The parity symmetry condition on the transfer imposes that if $\left|k\right\rangle$ is an eigenvector of $T$, then $\sigma_{x}\left|k\right\rangle$ is also an eigenvector of $k$. Importantly, if $T\left|k\right\rangle=\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0muk}\left|k\right\rangle$, then $T\sigma_{x}\left|k\right\rangle=T^{-1}\left|k\right\rangle=\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0muk}\left|k\right\rangle$. This lets us identify $\sigma_{x}\left|k\right\rangle=\left|-k\right\rangle$, and shows that the magnitude of the two entries in the transfer matrix are equal. That is, $\left|k\right\rangle=(1,\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu\varphi})$.
The converse also holds true: If $\left|k\right\rangle$ and $\sigma_{x}\left|k\right\rangle$ are eigenvectors with eigenvalue $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0muk}$ and $\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0muk}$. Then $T$ is parity-symmetric.
This lets us write any Bloch eigenvector in the position basis as
@@
Here $\mathcal{N}$ is a normalisation constant and the entries in the dots are not important (see Appendix A for details).
@@
$\phi(k)=\mathrm{arg}\left[\frac{\langle 0|k\rangle}{\langle 1|k\rangle}\left\langle k\right|T\left|k\right\rangle\right]=k-\varphi(k),$ (10)
where $\left|0\right\rangle$ and $\left|1\right\rangle$ are the standard position basis vectors, and $\varphi(k)$ is the relative phase of the two entries in the $T$-matrix eigenvector in this basis.
Figure 2: Bulk-edge correspondence for the Zak phase, as viewed in $\mathbb{R}$), and projected to $n$ given by Eq. (7). The paths $T(E)$ for 4-band SSH4 model is plotted for various $\gamma_{13}/\sqrt{\gamma_{2}\gamma_{4}}$ ratios. The green regions denote transfer matrices which support bulk modes (allowed bands regions), while the brown lines denote transfer matrices which support edge modes. The purple dashed boundaries denote the points where $\phi=\pi$, and the red dashed curves denote the points where $\phi=0$. The Zak phase $\phi_{\text{Zak}}$ is $\pi$ when the $T(E)$ curve enters and exits from boundaries of different colours, and it is 0 when the two colours are the same. As such, we see that the creation of edge states (magenta stars) is necessarily accompanied by the change in Zak phase of the two neighbouring bands.
@@ IV.1 Tetramer SSH4 Model
We see that when $\bar{\mu}=0$, the $T$ matrix may support a topological edge state with $T$-eigenvalue $\gamma_{3}\gamma_{1}/(\gamma_{4}\gamma_{2})$. That is, the SSH4 model seems to act just like 2 jagged copies of the dimer SSH model, where the suppression factor for the edge state is just the product of the suppression factors for each dimer. However, we note that, even if $\gamma_{1}>\gamma_{2}$, the SSH4 chain may support a topological edge state as long as $\gamma^{2}_{13}/\gamma_{2}<\gamma_{4}$. Furthermore, even if $\gamma_{4}$ is the weakest coupling, we can always choose $\gamma_{13},\gamma_{2}$ so that the eigenvalue remains less than 1, and an edge state is supported. This shows that the ‘limiting to different groupings’ argument frequently used in the discussion of the SSH model and the Kitaev chain (see Refs. [30, 31] and Fig. 3) is not so easily generalisable to more complicated systems. This is not the only topological edge state of the SSH4 model. The $T$ matrix also supports edge states when $\bar{\mu}=\pm\sqrt{\gamma_{13}^{2}+\gamma_{2}^{2}}$. In this case the suppression factor is $\gamma_{2}/\gamma_{4}$, and it looks like the edge state of an SSH chain with the $\gamma_{1}=\gamma_{3}$ couplings ignored.
@@
Computing the Zak phase of each band as in Fig. 4 shows that it is quantised. This verifies that our restricted SSH4 model is indeed topological. Our analysis also shows that the topological invariant of the SSH4 model is $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
Finally, our analysis of the SSH4 shows that unlike the Bloch Hamiltonian, the transfer matrix lends itself easily to expanding the size of the unit cell. This comes at the expense of making long-range couplings more difficult to introduce. However, this also means that the Hamiltonians of continuous systems can still be described by $2\times 2$ transfer matrices, and as such is susceptible to our analysis thus far. Our method therefore generalises to any real one-dimensional quantum mechanical or photonic systems, without discretisation.
@@ IV.2 Dimerised Photonic Crystal
Figure 5: (a) Dimerised dielectric photonic crystal inspired by the SSH model. We have assumed for simplicity that the material is nonmagnetic. In the coupled mode approximation, each high index layer (dark) is modelled as a waveguide and the low index layer serves only for separation. The unit cell is chosen so that a fraction $1-q$ of the low index layer is at the left edge. For $q=0.5$ the unit cell is symmetric. (b) Numerical $T$-matrix calculation of the band structure of the asymmetric DPC with $q=1$. The edge states shown as red lines. The dimensionless parameters used are as follows: $\varepsilon_{1}=4,\varepsilon_{2}=2,k_{0}h_{1}=k_{0}h_{2}=2.$ The host material parameters has been chosen to be $\varepsilon_{h}=\mu_{h}=1$. We see that there is an edge state below the two bands, which is not predicted by the SSH model. In addition, the edge state in between the two bands appears in the wrong side $\delta h<0$ compared to the SSH model. Since the unit cell is asymmetric (i.e. Eq. (7) is violated), the edge states are not topologically protected, and the corresponding plot in $\mathrm{SL}(2,\mathbb{R})$ space is going to be 3 dimensional.
We may compute the band structure of TM waves in the DPC with $q=1$ using the $T$-matrix numerically, with the bulk bands computed using Bloch’s theorem, and the edge states by numerically finding solutions to (23) in the following manner: The parameter space spanned by $\beta$ and $\delta h$ is discretised to a grid of size $500\times 50$ respectively. A cell is then marked as an edge state if (23) is satisfied within an adjustable tolerance. This results Fig. 5 (b) and we see that the results are unlike the SSH model. The edge state appears on the wrong side $\delta h<0$, which in the coupled mode approximation means $\gamma_{1}>\gamma_{2}$. In addition, we see an extended edge state appearing below the two bands, which spans through both regions $\delta h<0$ and $\delta h>0$. This was the original motivation of this work: to understand the nature of this extended edge state, and see if it is topologically protected.
@@
The condition that $T$ is parity invariant, Eq. (7), is exactly the condition that the unit cell is symmetric $q=1/2$. In this case, all edge states are topological and the $\mathrm{SL}(2,\mathbb{R})$ trajectories for the TM case are illustrated in Fig. 6. The TE case gives similar results, so we will omit it for brevity. From this figure, we see that edge states are supported if and only if $\delta h>0$, in agreement with the coupled mode approximation. There seems to be no extended edge states in this case. As a result, we see that the extended edge mode in Fig. 5 (b) is a result of the defect introduced by the asymmetric prescription of the unit cell, and is not a property of the bulk.
@@
We have characterised the edge states of the DPC for the allowed values of $\beta$ such that the field decays away from the periodic structure. The extended edge state is understood to arise from the defect introduced by the asymmetric perturbation $q\neq 0$, and its existence seems to be weakly tied to the topological edge state between the two SSH-like bands predicted by the coupled mode approximation.

### v2

@@ Abstract
We suggest and develop a novel approach for describing topological properties of a periodic system purely from the transfer matrix associated to a unit cell. This approach is employed to obtain a simple pictorial proof of the Zak phase bulk-edge correspondence for any one dimensional system. We apply the transfer matrix method to study some generalisations of the Su-Schrieffer-Heeger (SSH) model, such as the tetramer SSH4 model and a dimerised one-dimensional photonic crystal. We thus obtain simple conditions for the existence of topologically protected edge states in these models.
@@ I Introduction
Of these topological insulators, one of the simplest toy models is the Su-Schrieffer-Heeger (SSH) model, which was first introduced to model electron hopping in long-chain polyenes [6]. Despite its simplicity, the SSH model and its generalisations are often used in experiment to demonstrate techniques for detecting topological phase transitions [7], as well as to demonstrate practical applications of topological edge states [8]. Theoretically, the SSH model and its generalisations serve as a starting point to test new methods for analysing topological insulators [9, 10, 11]. Generalisations of the SSH model often also admit exciting nontrivial behaviour, which may shed light on the nature of topological phases [12], or present opportunities for technological applications [13]. On the other hand, due to its simplicity, many models in different fields reduce to the SSH model. For example, the SSH model has recently been realised in acoustic [14] and magnonic [15] systems.
In this article, we introduce a new method to analyse one-dimensional topological insulators using the transfer matrix. This transfer matrix method offers an approach complementary to the standard Bloch Hamiltonian method. As a preview, a system with $N\times N$ sites per unit cell and $k$-nearest neighbour hoppings would admit a Bloch Hamiltonian of size $N\times N$, while the transfer matrix would have a size of $2k\times 2k$. In addition, while the Bloch Hamiltonian works best with periodic boundary conditions, the transfer matrix method can be naturally equipped with open or closed boundary conditions. A consequence of this is that the edge states appear naturally within the transfer matrix method, so that we may characterise the edge states more explicitly. We thus apply the transfer matrix method to inversion-symmetric models and obtain simple conditions for the appearance of their topologically protected edge states.
The paper is organised as follows. In Sec. II, We begin by reviewing the transfer matrix method and apply it to solve the semi-infinite SSH model. In Sec. III, we demonstrate how this the transfer matrix approach can be employed to obtain topological invariants for the Hamiltonian. This method involves assigning to each Hamiltonian a path on a manifold with the endpoints fixed. We then show how our method connects to bulk-band topological invariants such as the Zak phase. In doing so, we obtain a short, but rigorous proof of the bulk-edge correspondence in general one dimensional systems possessing inversion symmetry. Finally, in Sec. IV we show how our method can be applied to more complicated models such as the generalized SSH model and a dimerised photonic crystal.
@@ II Transfer Matrix of the SSH model
The transfer ($T$) matrix is a powerful method of solving linear one-dimensional systems. It has previously also been considered in the study of topological insulators [16, 17, 18, 19, 11], though our approach is considerably different. To derive the transfer matrix, it is convenient to rewrite the time-independent Schrödinger equation $H\psi=E\psi$ in matrix form:
@@
where $\bm{a}_{i}=(a_{2i},a_{2i+1})^{T}$ and $a_{2i}$ is the rightmost amplitude of the previous unit cell and $a_{2i+1}$ is the first (leftmost) amplitude of the current unit cell. In this way, for a system of $N$ sites, we have introduced 2 more amplitudes in our transfer matrix description, $a_{0}$ and $a_{N+1}$. These will be important in the next paragraph when we consider boundary conditions. The gaussian elimination process also shows that in general, any one-dimensional periodic dynamical system admits a set of transfer matrices that is unique up to basis changes or symmetry transformations of the model [^1] This freedom is important as any hermitian tridiagonal Hamiltonian admits a gauge transformation to a real-symmetric tridiagonal Hamiltonian. This choice of basis will become important for the arguments in the later sections..
The main advantage of the $T$-matrix formalism appears when we consider semi-infinite and finite chains. For the SSH model, the Dirichlet boundary conditions as applied in Eq. (2) is equivalent to setting the amplitude $a_{0}$ to the left of the first amplitude to be zero. As a result, the amplitudes over the entire chain are determined by the amplitude at the first site by the $T$-matrix:
@@
The only cases where Bloch’s theorem still applies are the cases when both eigenvalues of $T$ are equal, and edge modes, where the left and right edge modes do not couple. So, the left edge mode is obtained by the choice of $E$ which makes $(0,1)^{T}$ an eigenvector of the $T$ matrix. This is easily read off from (5) to be $\mu_{1}=v_{1}-E=0$. Finally, substituting this into our $T$-matrix, we recover the well-known condition that the left edge mode exists if and only if $|\gamma_{1}/\gamma_{2}|<1$. A main advantage of this technique is that we are able to obtain the condition for an edge mode in the SSH model without any symmetry or hermiticity requirement on the Hamiltonian. As such, the condition $|\gamma_{1}/\gamma_{2}|<1$ works just as well for $\gamma_{1},\gamma_{2}\in\mathbb{C}$ (non-hermitian models, e.g. systems with gain, nonreciprocal hopping, etc), as well as for symmetry-broken SSH models with jagged on-site potentials.
@@ III Topological Protection in the Transfer Matrix Formalism
Figure 1: (a) Illustration of the SSH model. The open circles denote sites at which the electron can sit, and the lines denote the coupling strength between each site. In the SSH model, the couplings alternate between two values $\gamma_{1}$ and $\gamma_{2}$. (b) The band structure of the SSH model $(\mu_{1}=\mu_{2})$ embedded in $\mathrm{SL}(2,\mathbb{R})$ (3D Iwasawa parameters space). The parameter $k$ is only defined modulo $2\pi$. Each curve represents a Hamiltonian, or equivalently, a family of $T$-matrices $T(E)$ for a different value of $\gamma_{1}/\gamma_{2}$. When the curve touches one of the orange surfaces, an edge state is supported. The decay length of this edge state is given by $|\ln(a)|^{-1}$. The green regions are the allowed bands: Each point on the curve inside the green region corresponds to a bulk propagating mode.
@@ III.1 The $\mathbb{R}$) Picture for Topological Phases
$T=\begin{pmatrix}\cos(k)&\sin(k)\\ -\sin(k)&\cos(k)\end{pmatrix}\begin{pmatrix}\frac{1}{a}&\\ &a\end{pmatrix}\begin{pmatrix}1&\\ n&1\end{pmatrix}$ (6)
where $k,n\in\mathbb{R},a>0$ are the Iwasawa parameters, and they are smooth functions of the entries $T_{ij}$. By doing so, we can parameterise $T(E)$ as a curve in a three-dimensional space, where the coordinates are taken to be the Iwasawa parameters: $k(E)$, $a(E)$, $n(E)$. We plot a family of curves with various choices of $\gamma$ in Fig. 1 for the SSH model. Since $\det T\in\mathrm{SL}(2,\mathbb{R})$ the condition that the bulk states must stay bounded in the chain is automatically satisfied and we may identify the transfer matrices supporting bulk states by simply looking at the condition $\mathrm{Tr}(T)\leq 2$. On the other hand for a $T$ matrix to support an edge state, we require $(0,1)$ to be an eigenvector of $T$ with eigenvalue less than one. So, we need $k\in\pi\mathbb{Z}$ and $a<1$.
@@ III.2 Proof of the Zak Phase Bulk-Edge Correspondence
Here, we show how our formalism so far is connected to Bloch band topological invariants. In doing so, we will obtain a rigorous but pictorially simple proof of the bulk-edge correspondence. We stress that our results hold for any number of sites per unit cell, as well as for continuous quantum mechanical or photonic systems. This will become more clear in section IV.1 where we treat the case of $4$ sites per unit cell more explicitly. But for now, we will simply highlight a few key features when we increase the number of sites per unit cell.
For $N$ sites per unit cell, while the Bloch Hamiltonian grows to a size of $N\times N$, as long as interactions are kept to nearest-neighbours, the transfer matrix stays at a size of $2\times 2$. While the Bloch Hamiltonian acts on the amplitudes on the whole unit cell $(a_{jN+1},a_{jN+2},\cdots,a_{(j+1)N}),$ the transfer matrix acts only on the amplitudes $(a_{jN},a_{jN+1})$ to give $(a_{(j+1)N},a_{(j+1)N+1}).$
To connect between the transfer matrix picture and the Bloch Hamiltonian picture, let us first begin by analysing the Bloch eigenvector. Parity and time reversal [^3] The Hamiltonian considered in this paper has real entries, so it trivially satisfies time-reversal symmetry. This statemenet however is true for PT-symmetric Hamiltonians which are allowed to be complex. symmetry lets us write any Bloch eigenvector in the position basis as
@@
Here $\mathcal{N}$ is a normalisation constant, $\phi(K)$ is a real function, and the entries in the dots are not important (see Appendix A for details).
@@
$\phi(K)=\mathrm{arg}\left[\frac{\langle 0|K\rangle_{T}}{\langle 1|K\rangle_{T}}\left\langle K\right|_{T}T\left|K\right\rangle_{T}\right]=K-\varphi(K),$ (10)
where $\left|0\right\rangle$ and $\left|1\right\rangle$ are the standard position basis vectors [^4] For a finite chain of $N$ sites with Dirichlet boundary conditions this is defined to be a basis such that the boundary condition is given by $a_{0}=0$ and $a_{N+1}=0$ as in equation (5)., $\left|K\right\rangle_{T}$ is the eigenvector of $T$ with eigenvalue $\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0muK}$, and $\varphi(K)$ is the relative phase of the two entries in the $T$-matrix eigenvector in this basis. It should be noted here that both $\phi$ and $\varphi$ depend on the choice of basis. This is why the middle expression in (10) must be written in a basis-dependent way.
Figure 2: Bulk-edge correspondence for the Zak phase, as viewed in $\mathbb{R}$), and projected to $n$ given by Eq. (7). The paths $T(E)$ for 4-band SSH4 model is plotted for various $\gamma_{13}/\sqrt{\gamma_{2}\gamma_{4}}$ ratios. Note that the parameter $k$ is only defined modulo $2\pi$. The plot has been unfolded in much the same way as a band diagram can be unfolded beyond the first Brillouin zone. The green regions denote transfer matrices which support bulk modes (allowed bands regions), while the brown lines denote transfer matrices which support edge modes. The purple dashed boundaries denote the points where $\phi=\pi$, and the red dashed curves denote the points where $\phi=0$. The Zak phase $\phi_{\text{Zak}}$ is $\pi$ when the $T(E)$ curve enters and exits from boundaries of different colours, and it is 0 when the two colours are the same. As such, we see that the creation of edge states (magenta stars) is necessarily accompanied by the change in Zak phase of the two neighbouring bands.
@@ IV.1 Tetramer SSH4 Model
We see that when $\bar{\mu}=0$, the $T$ matrix may support a topological edge state with $T$-eigenvalue $\gamma_{3}\gamma_{1}/(\gamma_{4}\gamma_{2})$. That is, the SSH4 model seems to act just like 2 jagged copies of the dimer SSH model, where the $T$-eigenvalue for the edge state is just the product of the eigenvalues for each dimer. However, we note that, even if $\gamma_{1}>\gamma_{2}$, the SSH4 chain may support a topological edge state as long as $\gamma^{2}_{13}/\gamma_{2}<\gamma_{4}$. Furthermore, even if $\gamma_{4}$ is the weakest coupling, we can always choose $\gamma_{13},\gamma_{2}$ so that the eigenvalue remains less than 1, and an edge state is supported. This shows that the ‘limiting to different groupings’ argument frequently used in the discussion of the SSH model and the Kitaev chain (see Refs. [31, 32] and Fig. 3) is not so easily generalisable to more complicated systems. This is not the only topological edge state of the SSH4 model. The $T$ matrix also supports edge states when $\bar{\mu}=\pm\sqrt{\gamma_{13}^{2}+\gamma_{2}^{2}}$. In this case the eigenvalue is $\gamma_{2}/\gamma_{4}$, and it looks like the edge state of an SSH chain with the $\gamma_{1}=\gamma_{3}$ couplings ignored.
@@
Computing the Zak phase of each band as in Fig. 4 shows that it is quantised. This verifies that our restricted SSH4 model is indeed topologically nontrivial. Our analysis also shows that the group of topological invariants for the SSH4 model is $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
So far we have only analysed the SSH4 model protected by parity symmetry. In the dimer case, both parity symmetry and chiral symmetry impose the same requirement $v_{1}=v_{2}$, and both are individually responsible for the topological protection of the edge mode [25]. However, in the tetramer SSH4 case chiral symmetry imposes a different condition $v_{i}=0$ than parity symmetry. For a brief analysis, on the transfer matrix level, chiral symmetry is satisfied as long as
$\sigma_{z}T(E)\sigma_{z}=T(-E).$
This relation provides protection only to the edge mode between the second and third bands at $E=0$. The other edge modes protected by parity symmetry are no longer protected when only chiral symmetry is present.
Finally, our analysis of the SSH4 model shows that unlike the Bloch Hamiltonian, the transfer matrix lends itself easily to expanding the size of the unit cell. This comes at the expense of making long-range couplings more difficult to introduce. However, this also means that the Hamiltonians of continuous systems can still be described by $2\times 2$ transfer matrices, and as such is susceptible to our analysis thus far. Our method therefore generalises to any real one-dimensional quantum mechanical or photonic systems, without discretisation.
@@ IV.2 Dimerised Photonic Crystal
Figure 5: (a) Dimerised dielectric photonic crystal inspired by the SSH model. We have assumed for simplicity that the material is nonmagnetic. The unit cell is chosen so that a fraction $1-q$ of the low index layer is at the left edge. For $q=0.5$ the unit cell is symmetric. (b) Numerical $T$-matrix calculation of the band structure of the asymmetric DPC with $q=1$. The edge states shown as red lines. The dimensionless parameters used are as follows: $\varepsilon_{1}=4,\varepsilon_{2}=2,k_{0}h_{1}=k_{0}h_{2}=2.$ The host material parameters has been chosen to be $\varepsilon_{h}=\mu_{h}=1$. We see that there is an edge state below the two bands, which is not predicted by the SSH model. In addition, the edge state in between the two bands appears in the wrong side $\delta h<0$ compared to the SSH model. Since the unit cell is asymmetric (i.e. Eq. (7) is violated), the edge states are not topologically protected, and the corresponding plot in $\mathrm{SL}(2,\mathbb{R})$ space is going to be 3 dimensional.
@@
We may compute the band structure of TM waves in the DPC with $q=1$ using the $T$-matrix numerically, with the bulk bands computed using Bloch’s theorem, and the edge states by numerically finding solutions to (23) in the following manner: The parameter space spanned by $\beta$ and $\delta h$ is discretised to a grid of size $500\times 50$ respectively. A cell is then marked as an edge state if (23) is satisfied within an adjustable tolerance. This results Fig. 5 (b) and we see that the results are unlike the SSH model. The edge state appears on the wrong side $\delta h<0$, which in the coupled mode approximation (Appendix B) means $\gamma_{1}>\gamma_{2}$. In addition, we see an extended edge state appearing below the two bands, which spans through both regions $\delta h<0$ and $\delta h>0$.
The condition that $T$ is parity invariant, Eq. (7), is exactly the condition that the unit cell is symmetric $q=1/2$. In this case, all edge states are topological and the $\mathrm{SL}(2,\mathbb{R})$ trajectories for the TM case are illustrated in Fig. 6. The TE case gives similar results, so we will omit it for brevity. From this figure, we see that edge states are supported if and only if $\delta h>0$, in agreement with the coupled mode approximation in Appendix B. There seems to be no extended edge states in this case. As a result, we see that the extended edge mode in Fig. 5 (b) is a result of the defect introduced by the asymmetric prescription of the unit cell, and is not a property of the bulk.
@@
We have characterised the edge states of the DPC for the allowed values of $\beta$ such that the field decays away from the periodic structure. The extended edge state is understood to arise from the defect introduced by the asymmetric perturbation $q\neq 0$, and its existence seems to be weakly tied to the topological edge state between the two SSH-like bands predicted by the coupled mode approximation (Appendix B).